A sphere or radius moving with velocity through a fluid experiencces a retarding Force given by f=6ry vz nx where n is the viscosity of the liquid.use method of dimension to find value of x,y,z
1
Expert's answer
2021-08-23T14:12:13-0400

Gives

"F=6\pi r\eta v_t"

"F=6r^yv^zn^x"

Dimensions

"F=[M^1L^1T^{-2}]"

"r=[L]\\v=[LT^{-1}]"

"\eta=[M^1L^{-1}T^{-1}]"


"[M^1L^{-1}T^{-2}]=6[L]^x[LT^{-1}]^y[M^1L^{-1}T^{-1}]^z\\x=1,y=1,z=1"